A population consists of six scores: 11, 13, 15, 16, 17, and 19. If

a sampling distribution of size 3 is drawn from the population, what is the standard

error of the mean of the sampling distribution? complete solution


1
Expert's answer
2021-05-07T12:29:11-0400

Population consists of the numbers 11, 13, 15, 16, 17, and 19

Sampling distribution of size 3 can be given as




Mean of sampling distribution is "= \dfrac{303.27}{20} = 15.16"


Then, Standard Error can be calculated using the formula,

"SE = \dfrac{s}{\sqrt{n}}"


s = standard deviation of sample mean.

n = 20.